# Indirectly Parameterized Concrete Autoencoders

### v1

@@ 1 Introduction
Feature selection enables more parsimonious and interpretable models, and is therefore a fundamental task in machine learning and statistics in general and essential in several applications such as bioinformatics e.g., gene subset selection, neuroscience e.g., fMRI analysis, and fluid mechanics e.g., optimal sensor placement. Unfortunately, finding the optimal set of features is NP-hard (Amaldi & Kann, 1998).
@@ 2 Method
In this section we describe vanilla CAE, discuss its shortcomings and present our proposed improvements to CAE training. We introduce a new way to indirectly parameterize the Gumbel-Softmax distributions in CAE, which we call Indirect Parametrization (IP). We further propose a baseline diversity-encouraging regularization method using the Generalized Jensen Shannon Divergence (GJSD).
@@ 2.1 Concrete Autoencoders (CAE)
The CAE (Balın et al., 2019) architecture is competitive with state-of-the-art methods in neural network-based embedded feature selection. CAEs consists of two components: a concrete selection layer that performs differentiable feature selection on the input features (encoder) and an arbitrary neural network (decoder). A predictive network is used in place of the decoder for classification or regression tasks.
@@ 2.2 Challenges of CAE
We empirically demonstrate instability during training of CAEs (Figure 1, top). We interestingly show that this instability strongly correlates with the unique percentage, consistently across tasks and datasets (Figure 1, bottom).
@@
In the next subsection we describe our proposed IP-CAE which is later shown to alleviate the aforementioned shortcoming of vanilla CAE including instability, unique percentange, variance (Figure 5), training time (Table 3), and non-linear decoder (Figure 5(b)) leading to state-of-the-art performance for feature selection for both reconstruction (Table 1) and classification (Table 2) on all datasets considered.
@@ 2.4 A Comparison Between CAE and IP-CAE
For simplicity, we consider IP-CAE with $P=D$ and without the bias term, and thus have $\log{\boldsymbol{\alpha}}_{i}=\boldsymbol{W}\boldsymbol{\psi}_{i}$, with the following update rule (detailed derivations in Appendix C):
@@
where $\nabla\mathcal{L}$ is the gradient of the loss with respect to $\log\boldsymbol{\alpha}_{i}^{(t)}=\boldsymbol{W}^{(t)}\boldsymbol{\psi}_{i}^{(t)}=\boldsymbol{W}\boldsymbol{\psi}_{i}$, and the step-dependent matrix $\boldsymbol{T}_{i}\in\mathbb{R}^{D\times D}$ represents a learned transformation of the gradients. The transform affects the gradients in two ways: a linear transformation represented by $\boldsymbol{W}\boldsymbol{W}^{T}$ that is shared for all $i$, and a scaling by $\boldsymbol{\psi}_{i}^{T}(\boldsymbol{\psi}_{i}-\eta\boldsymbol{W}^{T}\nabla\mathcal{L})\in\mathbb{R}$, which is the dot product between $\boldsymbol{\psi}_{i}^{(t)}$ and $\boldsymbol{\psi}_{i}^{(t+1)}$. A geometric interpretation of the dot product is $||\boldsymbol{\psi}_{i}^{(t)}||_{2}||\boldsymbol{\psi}_{i}^{(t+1)}||_{2}\cos(\theta_{i})$ with $\theta_{i}$ being the angle between the two vectors. Inuitively, early in training when $\boldsymbol{\psi}_{i}$ is close to the origin, small changes in $\boldsymbol{\psi}_{i}$ can lead to large $\theta_{i}$, making the transform be dominated by $\boldsymbol{W}\boldsymbol{W}^{T}$. However, once a stable direction for $\boldsymbol{\psi}_{i}$ has been found such that $\theta_{i}$ is small, the scalar scaling part of the transformation also comes into play, effectively acting as a learned learning rate increase for $\log\boldsymbol{\alpha}_{i}$.
@@ 3 Related Work
Feature Selection using STGs Yamada et al. (2020), similarly to CAE, use stochastic nodes to sample features during training to perform differentiable joint optimization of feature selection and model. While CAE models the selection using $K$ categorical nodes in the concrete selector layer, Stochastic Gaussian Gates (STGs) uses $D$ Bernoulli nodes, each for one input feature, where $K$ denotes the desired number of optimal features, and $D$ denotes the total number of input features. The authors propose a novel reparametrization for Bernoulli variables using thresholded Gaussian variables to allow for differentiable learning. The use of Bernoulli gates is closely related to the Bernoulli-Gaussian model for linear regression with feature selection and $\ell_{0}$ regularization.
LassoNet (Lemhadri et al., 2021) is an extension of the popular Lasso method for regression. Although the classic Lasso has been efficient and useful for embedded feature selection in linear models, it is challenging to generalize it to neural network models (Cui & Wang, 2016). The LassoNet architecture achieves this by introducing residual connections from the input layer to the output of the network and applying an $\ell_{1}$ penalty term to that layer. The design principle is such that it allows for a feature to be selected by the model if and only if it also gets selected by the residual layer. The authors model this principle as an explicit constraint in the optimization problem and propose a projected proximal gradient optimization algorithm to ensure the constraint satisfiability during the process.
@@ 4 Experiments
Model
MNIST
MNIST-Fashion
ISOLET
COIL-20
Smartphone HAR
Mice Protein
STG
2.42E-02 $\pm$ 2.70E-06
1.80E-02 $\pm$ 2.84E-06
7.00E-03 $\pm$ 1.96E-06
5.00E-03 $\pm$ 6.75E-06
4.00E-03 $\pm$ 3.04E-06
1.17E-02 $\pm$ 6.43E-05
LassoNet
1.98E-02 $\pm$ 9.02E-06
1.96E-02 $\pm$ 4.23E-06
7.40E-03 $\pm$ 4.25E-06
6.24E-03 $\pm$ 5.42E-06
4.01E-03 $\pm$ 1.29E-06
1.16E-02 $\pm$ 4.45E-05
CAE
1.48E-02 $\pm$ 1.31E-04
1.36E-02 $\pm$ 1.17E-04
5.42E-03 $\pm$ 6.25E-05
4.21E-03 $\pm$ 7.69E-05
3.03E-03 $\pm$ 6.45E-05
8.71E-03 $\pm$ 2.61E-04
GJSD
1.45E-02 $\pm$ 1.14E-04
1.23E-02 $\pm$ 9.26E-05
4.86E-03 $\pm$ 4.52E-05
2.80E-03 $\pm$ 2.85E-05
2.57E-03 $\pm$ 3.50E-05
8.23E-03 $\pm$ 2.99E-04
IP-CAE
1.36E-02 $\pm$ 5.35E-05
1.17E-02 $\pm$ 3.15E-05
4.38E-03 $\pm$ 1.28E-05
2.48E-03 $\pm$ 1.39E-05
2.03E-03 $\pm$ 2.75E-05
6.90E-03 $\pm$ 1.25E-04
Model
MNIST
MNIST-Fashion
ISOLET
COIL-20
Smartphone HAR
Mice Protein
STG
$92.29\;\;{\color[rgb]{0.7,0.7,0.7}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0.7,0.7}\pgfsys@color@gray@stroke{0.7}\pgfsys@color@gray@fill{0.7}\pm 0.30}$
$80.85\;\;{\color[rgb]{0.7,0.7,0.7}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0.7,0.7}\pgfsys@color@gray@stroke{0.7}\pgfsys@color@gray@fill{0.7}\pm 0.27}$
$84.95\;\;{\color[rgb]{0.7,0.7,0.7}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0.7,0.7}\pgfsys@color@gray@stroke{0.7}\pgfsys@color@gray@fill{0.7}\pm 0.31}$
$96.80\;\;{\color[rgb]{0.7,0.7,0.7}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0.7,0.7}\pgfsys@color@gray@stroke{0.7}\pgfsys@color@gray@fill{0.7}\pm 0.25}$
$88.80\;\;{\color[rgb]{0.7,0.7,0.7}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0.7,0.7}\pgfsys@color@gray@stroke{0.7}\pgfsys@color@gray@fill{0.7}\pm 0.08}$
$68.24\;\;{\color[rgb]{0.7,0.7,0.7}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0.7,0.7}\pgfsys@color@gray@stroke{0.7}\pgfsys@color@gray@fill{0.7}\pm 1.11}$
LassoNet
$90.06\;\;{\color[rgb]{0.7,0.7,0.7}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0.7,0.7}\pgfsys@color@gray@stroke{0.7}\pgfsys@color@gray@fill{0.7}\pm 0.33}$
$78.28\;\;{\color[rgb]{0.7,0.7,0.7}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0.7,0.7}\pgfsys@color@gray@stroke{0.7}\pgfsys@color@gray@fill{0.7}\pm 0.36}$
$84.33\;\;{\color[rgb]{0.7,0.7,0.7}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0.7,0.7}\pgfsys@color@gray@stroke{0.7}\pgfsys@color@gray@fill{0.7}\pm 0.28}$
$89.37\;\;{\color[rgb]{0.7,0.7,0.7}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0.7,0.7}\pgfsys@color@gray@stroke{0.7}\pgfsys@color@gray@fill{0.7}\pm 0.47}$
$92.44\;\;{\color[rgb]{0.7,0.7,0.7}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0.7,0.7}\pgfsys@color@gray@stroke{0.7}\pgfsys@color@gray@fill{0.7}\pm 0.11}$
$77.12\;\;{\color[rgb]{0.7,0.7,0.7}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0.7,0.7}\pgfsys@color@gray@stroke{0.7}\pgfsys@color@gray@fill{0.7}\pm 0.80}$
CAE
$83.10\;\;{\color[rgb]{0.7,0.7,0.7}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0.7,0.7}\pgfsys@color@gray@stroke{0.7}\pgfsys@color@gray@fill{0.7}\pm 1.23}$
$73.19\;\;{\color[rgb]{0.7,0.7,0.7}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0.7,0.7}\pgfsys@color@gray@stroke{0.7}\pgfsys@color@gray@fill{0.7}\pm 0.82}$
$75.82\;\;{\color[rgb]{0.7,0.7,0.7}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0.7,0.7}\pgfsys@color@gray@stroke{0.7}\pgfsys@color@gray@fill{0.7}\pm 2.31}$
$80.70\;\;{\color[rgb]{0.7,0.7,0.7}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0.7,0.7}\pgfsys@color@gray@stroke{0.7}\pgfsys@color@gray@fill{0.7}\pm 2.93}$
$82.72\;\;{\color[rgb]{0.7,0.7,0.7}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0.7,0.7}\pgfsys@color@gray@stroke{0.7}\pgfsys@color@gray@fill{0.7}\pm 0.80}$
$63.10\;\;{\color[rgb]{0.7,0.7,0.7}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0.7,0.7}\pgfsys@color@gray@stroke{0.7}\pgfsys@color@gray@fill{0.7}\pm 6.51}$
GJSD
$84.38\;\;{\color[rgb]{0.7,0.7,0.7}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0.7,0.7}\pgfsys@color@gray@stroke{0.7}\pgfsys@color@gray@fill{0.7}\pm 1.50}$
$74.13\;\;{\color[rgb]{0.7,0.7,0.7}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0.7,0.7}\pgfsys@color@gray@stroke{0.7}\pgfsys@color@gray@fill{0.7}\pm 0.64}$
$77.56\;\;{\color[rgb]{0.7,0.7,0.7}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0.7,0.7}\pgfsys@color@gray@stroke{0.7}\pgfsys@color@gray@fill{0.7}\pm 0.82}$
$82.10\;\;{\color[rgb]{0.7,0.7,0.7}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0.7,0.7}\pgfsys@color@gray@stroke{0.7}\pgfsys@color@gray@fill{0.7}\pm 3.72}$
$84.78\;\;{\color[rgb]{0.7,0.7,0.7}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0.7,0.7}\pgfsys@color@gray@stroke{0.7}\pgfsys@color@gray@fill{0.7}\pm 1.04}$
$68.43\;\;{\color[rgb]{0.7,0.7,0.7}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0.7,0.7}\pgfsys@color@gray@stroke{0.7}\pgfsys@color@gray@fill{0.7}\pm 7.75}$
IP-CAE
94.07  $\pm$0.37 82.68  $\pm$0.80 91.85  $\pm$0.55 97.92  $\pm$0.57 93.71  $\pm$0.62 94.26   $\pm$ 1.48
In this section, we evaluate our proposed method on several datasets. Table 4 in Appendix A provides an overview of the datasets used. Across all experiments, we perform an ablation of the standard CAE, only IP, only the GJSD term, and both IP and the GJSD term. The hyperparameter P in $\Psi=\mathbb{R}^{K\times P}$ controlling the embedding dimension of the IP selector layer, is not tuned in our experiments. Instead, we simply set $P=D$, where $D$ is the number of features for each dataset. The regularization strength hyperparameter for GJSD, $\lambda$, was tuned in $\{0,0.0005,0.005,0.05\}$. Following CAE, we use a fixed learning rate of 0.001 with the Adam optimizer with moving average coefficients $\boldsymbol{\beta}=(0.9,0.999)$ and no weight decay, for all experiments and datasets. We train every model for 200 epochs, and select the weights corresponding to the best validation loss for test set evaluation. In all experiments, unless otherwise specified, we use an MLP with one hidden layer of 200 nodes for the decoder network. For the hidden activation, we use LeakyReLU with a slope of 0.2. For all experiments, we perform 10 repetitions and report the mean quantity. Any confidence intervals correspond to one standard deviation.
MNIST and MNIST-Fashion (Deng, 2012; Xiao et al., 2017) consist of $28\times 28$ greyscale images depicting digits and clothing items respectively. The supervised classification task is to predict the item from the pixel values.
ISOLET (Cole & Fanty, 1994) consists of preprocessed speech data of test subjects speaking all 52 letters of the English alphabet. The supervised classification task is to predict the spoken letter from the 617-dimensional speech data.
@@
Mice Protein Expression (Higuera et al., 2015a) consists of protein expressions from two groups of mice; control and trisomic mice. The supervised classification task is to predict the label consisting of the group, stimulation, and treatment of the mice. 7 proteins have missing values for one or more mice and these values were imputed with the average protein expression level of examples belonging to the same class of mice, following the dataset authors(Higuera et al., 2015b).

### v2

@@ 1 Introduction
Feature selection is a fundamental task in machine learning and statistics, enabling more parsimonious and interpretable models. It is essential in several applications such as bioinformatics e.g., gene subset selection, neuroscience e.g., fMRI analysis, and fluid mechanics e.g., optimal sensor placement. Moreover, feature selection is often used for regularization. Unfortunately, finding the optimal selection is NP-hard (Amaldi & Kann, 1998).
@@ 2 Method
In this section we describe the vanilla CAE, discuss its shortcomings and present our proposed improvements to CAE training. We introduce a new way to indirectly parameterize the Gumbel-Softmax distributions in CAE, which we call Indirect Parametrization (IP). We further propose a baseline diversity-encouraging regularization method using the Generalized Jensen Shannon Divergence (GJSD).
@@ 2.1 Concrete Autoencoders (CAE)
The CAE (Balın et al., 2019) architecture is competitive with state-of-the-art methods in neural network-based embedded feature selection. CAEs consist of two components: a concrete selection layer that performs differentiable feature selection on the input features (encoder) and an arbitrary neural network (decoder). A predictive network is used in place of the decoder for classification or regression tasks.
@@ 2.2 Challenges of CAE
We empirically demonstrate instability during training of CAEs (Figure 1, top). Interestingly, our results show that this instability strongly correlates with the unique percentage, consistently across tasks and datasets (Figure 1, bottom).
@@
In the next subsection we describe our proposed IP-CAE which is later shown to alleviate the aforementioned shortcoming of vanilla CAE including instability, unique percentange, variance (Figure 5), training time (Table 3), and non-linear decoder (Figure 6) leading to state-of-the-art performance for feature selection for both reconstruction (Table 1) and classification (Table 2) on all datasets considered.
@@ 2.4 A Comparison Between CAE and IP-CAE
For simplicity, we consider IP-CAE with $P=D$ and without the bias term, and thus have $\log{\boldsymbol{\alpha}}_{i}=\boldsymbol{W}\boldsymbol{\psi}_{i}$, with the following update rule (detailed derivations in Appendix D):
@@
where $\nabla\mathcal{L}$ is the gradient of the loss with respect to $\log\boldsymbol{\alpha}_{i}^{(t)}=\boldsymbol{W}^{(t)}\boldsymbol{\psi}_{i}^{(t)}=\boldsymbol{W}\boldsymbol{\psi}_{i}$, and the step-dependent matrix $\boldsymbol{T}_{i}\in\mathbb{R}^{D\times D}$ represents a learned transformation of the gradients. The transform affects the gradients in two ways: a linear transformation represented by $\boldsymbol{W}\boldsymbol{W}^{T}$ that is shared for all $i$, and a scaling by $\boldsymbol{\psi}_{i}^{T}(\boldsymbol{\psi}_{i}-\eta\boldsymbol{W}^{T}\nabla\mathcal{L})\in\mathbb{R}$, which is the dot product between $\boldsymbol{\psi}_{i}^{(t)}$ and $\boldsymbol{\psi}_{i}^{(t+1)}$. A geometric interpretation of the dot product is $||\boldsymbol{\psi}_{i}^{(t)}||_{2}||\boldsymbol{\psi}_{i}^{(t+1)}||_{2}\cos(\theta_{i})$ with $\theta_{i}$ being the angle between the two vectors. Empirically, we have found that the learned rescaling changes throughout training, and our results suggest that these changes are beneficial. Because of the interactions between $\boldsymbol{W}$, $\phi$, and $\nabla\mathcal{L}$, the effect throughout training may be elaborate. We include additional experiments exploring the update’s behavior in Appendix C.
@@ 3 Related Work
Feature selection using STGs
Yamada et al. (2020), similarly to CAE, use stochastic nodes to sample features during training to perform differentiable joint optimization of feature selection and model. While CAE models the selection using $K$ categorical nodes in the concrete selector layer, Stochastic Gaussian Gates (STGs) uses $D$ Bernoulli nodes, each for one input feature, where $K$ denotes the desired number of optimal features, and $D$ denotes the total number of input features. The authors propose a novel reparametrization for Bernoulli variables using thresholded Gaussian variables to allow for differentiable learning. The use of Bernoulli gates is closely related to the Bernoulli-Gaussian model for linear regression with feature selection and $\ell_{0}$ regularization.
LassoNet
Lemhadri et al. (2021) extend the popular Lasso (Tibshirani, 1996) method for regression. Although the classic Lasso has been efficient and useful for embedded feature selection in linear models, it is challenging to generalize it to neural network models (Cui & Wang, 2016). The LassoNet architecture achieves this by introducing residual connections from the input layer to the output of the network and applying an $\ell_{1}$ penalty term to that layer. The design principle is such that it allows for a feature to be selected by the model if and only if it also gets selected by the residual layer. The authors model this principle as an explicit constraint in the optimization problem and propose a projected proximal gradient optimization algorithm to ensure the constraint satisfiability during the process.
Model MNIST MNIST-Fashion ISOLET COIL-20 Smartphone HAR Mice Protein
STG 2.42E-02 $\pm$ 2.70E-06 1.80E-02 $\pm$ 2.84E-06 7.00E-03 $\pm$ 1.96E-06 5.00E-03 $\pm$ 6.75E-06 4.00E-03 $\pm$ 3.04E-06 1.17E-02 $\pm$ 6.43E-05
LassoNet 1.98E-02 $\pm$ 9.02E-06 1.96E-02 $\pm$ 4.23E-06 7.40E-03 $\pm$ 4.25E-06 6.24E-03 $\pm$ 5.42E-06 4.01E-03 $\pm$ 1.29E-06 1.16E-02 $\pm$ 4.45E-05
CAE 1.48E-02 $\pm$ 1.31E-04 1.36E-02 $\pm$ 1.17E-04 5.42E-03 $\pm$ 6.25E-05 4.21E-03 $\pm$ 7.69E-05 3.03E-03 $\pm$ 6.45E-05 8.71E-03 $\pm$ 2.61E-04
GJSD 1.45E-02 $\pm$ 1.14E-04 1.23E-02 $\pm$ 9.26E-05 4.86E-03 $\pm$ 4.52E-05 2.80E-03 $\pm$ 2.85E-05 2.57E-03 $\pm$ 3.50E-05 8.23E-03 $\pm$ 2.99E-04
IP-CAE 1.36E-02 $\pm$ 5.35E-05 1.17E-02 $\pm$ 3.15E-05 4.38E-03 $\pm$ 1.28E-05 2.48E-03 $\pm$ 1.39E-05 2.03E-03 $\pm$ 2.75E-05 6.90E-03 $\pm$ 1.25E-04
Model MNIST MNIST-Fashion ISOLET COIL-20 Smartphone HAR Mice Protein
STG $92.29\;\;{\color[rgb]{0.7,0.7,0.7}\pm 0.30}$ $80.85\;\;{\color[rgb]{0.7,0.7,0.7}\pm 0.27}$ $84.95\;\;{\color[rgb]{0.7,0.7,0.7}\pm 0.31}$ $96.80\;\;{\color[rgb]{0.7,0.7,0.7}\pm 0.25}$ $88.80\;\;{\color[rgb]{0.7,0.7,0.7}\pm 0.08}$ $68.24\;\;{\color[rgb]{0.7,0.7,0.7}\pm 1.11}$
LassoNet $90.06\;\;{\color[rgb]{0.7,0.7,0.7}\pm 0.33}$ $78.28\;\;{\color[rgb]{0.7,0.7,0.7}\pm 0.36}$ $84.33\;\;{\color[rgb]{0.7,0.7,0.7}\pm 0.28}$ $89.37\;\;{\color[rgb]{0.7,0.7,0.7}\pm 0.47}$ $92.44\;\;{\color[rgb]{0.7,0.7,0.7}\pm 0.11}$ $77.12\;\;{\color[rgb]{0.7,0.7,0.7}\pm 0.80}$
CAE $83.10\;\;{\color[rgb]{0.7,0.7,0.7}\pm 1.23}$ $73.19\;\;{\color[rgb]{0.7,0.7,0.7}\pm 0.82}$ $75.82\;\;{\color[rgb]{0.7,0.7,0.7}\pm 2.31}$ $80.70\;\;{\color[rgb]{0.7,0.7,0.7}\pm 2.93}$ $82.72\;\;{\color[rgb]{0.7,0.7,0.7}\pm 0.80}$ $63.10\;\;{\color[rgb]{0.7,0.7,0.7}\pm 6.51}$
GJSD $84.38\;\;{\color[rgb]{0.7,0.7,0.7}\pm 1.50}$ $74.13\;\;{\color[rgb]{0.7,0.7,0.7}\pm 0.64}$ $77.56\;\;{\color[rgb]{0.7,0.7,0.7}\pm 0.82}$ $82.10\;\;{\color[rgb]{0.7,0.7,0.7}\pm 3.72}$ $84.78\;\;{\color[rgb]{0.7,0.7,0.7}\pm 1.04}$ $68.43\;\;{\color[rgb]{0.7,0.7,0.7}\pm 7.75}$
IP-CAE 94.07  $\pm$0.37 82.68  $\pm$0.80 91.85  $\pm$0.55 97.92  $\pm$0.57 93.71  $\pm$0.62 94.26   $\pm$ 1.48
@@
Other feature selection methods
Deep Lasso (Cherepanova et al., 2023) is another generalization of the Lasso, different from LassoNet. Instead of penalizing the weights directly, the authors suggest penalizing the gradient and recovering the classic Lasso as a special case of their method. Another approach is training sparse neural networks where the sparse input layer naturally performs feature selection (Louizos et al., 2018; Sokar et al., 2022). Although not trivially extended to neural networks, sparse priors formalize and extend the Lasso approach (Carvalho et al., 2009; Ročková & George, 2018). Deep Knockoffs (Romano et al., 2020) use deep generative models to enhance knockoff machines (Barber & Candès, 2015), a powerful method for statistical variable selection. Selecting measurements in compressed sensing, an important problem in medical imaging, has seen the development of similar methods to those in feature selection (Bakker et al., 2020; Huijben et al., 2019). Finally, instance-wise feature selection extends the problem of global feature selection to predict selections per sample, providing a hard, thresholded explanation of the network’s prediction (Yoon et al., 2018; Chen et al., 2018).
@@ 4 Experiments
In this section, we evaluate our proposed method on several datasets. Table 4 in Appendix A provides an overview of the datasets used. Across all experiments, we perform an ablation of the standard CAE, only IP, only the GJSD term, and both IP and the GJSD term. The hyperparameter $P$ in $\Psi=\mathbb{R}^{K\times P}$ controlling the embedding dimension of the IP selector layer, is not tuned in our experiments. Instead, we simply set $P=D$, where $D$ is the number of features for each dataset. The regularization strength hyperparameter for GJSD, $\lambda$, was tuned in $\{0,0.0005,0.005,0.05\}$. Following CAE, we use a fixed learning rate of 0.001 with the Adam optimizer with moving-average coefficients $\boldsymbol{\beta}=(0.9,0.999)$ and no weight decay, for all experiments and datasets. We train every model for 200 epochs, and select the weights corresponding to the best validation loss for test set evaluation. In all experiments, unless otherwise specified, we use an MLP with one hidden layer of 200 nodes for the decoder network. For the hidden activation, we use LeakyReLU with a slope of 0.2. For all experiments, we perform 10 repetitions and report the mean quantity. Any confidence intervals correspond to one standard deviation.
MNIST and MNIST-Fashion (LeCun et al., 1998; Xiao et al., 2017) consist of $28\times 28$ greyscale images depicting digits and clothing items respectively. The supervised classification task is to predict the item from the pixel values.
ISOLET (Fanty & Cole, 1990) consists of preprocessed speech data of test subjects speaking all 52 letters of the English alphabet. The supervised classification task is to predict the spoken letter from the 617-dimensional speech data.
@@
Mice Protein Expression (Higuera et al., 2015) consists of protein expressions from two groups of mice; control and trisomic mice. The supervised classification task is to predict the label consisting of the group, stimulation, and treatment of the mice. 7 proteins have missing values for one or more mice and these values were imputed with the average protein expression level of examples belonging to the same class of mice, following the dataset authors.